OpenStudy (anonymous):

find the seventh term of the geometric sequence with a3=125 and r=5

OpenStudy (anonymous):

geometric sequence \[a_n = a + r^{n-1}\] \[a_3 = 125\] \[r = 5\] \[125 = a + 5^{3-1}\] find a then \[a_7 = a+r^{7-1}\]